# The separating Noether number of small groups

### v1

@@ 6.1. The group $\mathrm{Dic}_{8}\times\mathrm{C}_{2}$
(note that $\langle a^{2}\rangle$ is the commutator subgroup of $G$, so $a^{2}$ is in the kernel of any $1$-dimensional representation of $G$). For $j=1,2,3,4$ denote by $W_{j}$ the vector space $K^{2}$ endowed with the representation $\psi_{j}$, and for $\chi\in\widehat{G}$ denote by $U_{\chi}$ the vector space $K$ endowed with the representation $\chi$, and set
$U:=\bigoplus_{\chi\in\widehat{G}}U_{\chi}.$
@@ 7. The group $\mathrm{S}_{3}\times\mathrm{C}_{3}$
For example, $q_{xx}=x_{1}x_{2}$, $q_{yz}=\frac{1}{2}(y_{1}z_{2}+y_{2}z_{1})$, $p_{xxx}=x_{1}^{3}+x_{2}^{3}$, $p_{xxy}=x_{1}^{2}y_{1}+x_{2}^{2}y_{2}$. It is well known (see for example [32, Theorem 4.1]) that the elements $q_{\xi\eta}$, $p_{\xi\eta\zeta}$ minimally generate $K[W_{1}\oplus W_{2}\oplus W_{3}]^{\mathrm{S}_{3}}$, where $\mathrm{S}_{3}$ is identified with the subgroup $\langle a,b\rangle$ of $G$. Set
@@
Assume next that $|\mathrm{Stab}_{G}(w)|=2$; this means that $H:=\mathrm{Stab}_{G}(w)$ is $\langle b\rangle$, $\langle ab\rangle$, or $\langle a^{2}b\rangle$, and $q_{yy}(w)\neq 0$. For any $H$-invariant monomial $T:=t_{\chi}\cdots t_{\chi^{\prime}}$ with $\deg(T)\leq 2$, either $T$, $q_{yy}T$, or $q_{yy}^{2}T$ is a $G$-invariant of degree at most $6$. It follows that $T(u)=T(u^{\prime})$ for any $H$-invariant of degree at most $2$. Therefore there exists an $h\in H$ with $h\cdot u=u^{\prime}$ (note that $\beta_{\mathrm{sep}}^{K}(H)=\beta^{K}(H)=2$).
Assume next that $|\mathrm{Stab}_{G}(w)|=3$; this means that $H:=\mathrm{Stab}_{G}(w)$ is $\langle ac\rangle$ or $\langle ac^{2}\rangle$, and $p^{-}_{yyy}(w)\neq 0$. For any $H$-invariant monomial $T:=t_{\chi}\cdots t_{\chi^{\prime}}$ with $\deg(T)\leq 3$, either $T$ or $p^{-}_{yyy}T$ is a $G$-invariant of degree at most $6$. It follows that $T(u)=T(u^{\prime})$ for any $H$-invariant of degree at most $3$. Therefore there exists an $h\in H$ with $h\cdot u=u^{\prime}$ (note that $\beta_{\mathrm{sep}}^{K}(H)=\beta^{K}(H)=3$).
@@ 8.3. The group $\mathrm{C}_{7}\rtimes\mathrm{C}_{3}$
It is proved in [9, Theorem 4.1] that for any prime $p$ congruent to $1$ modulo $3$, the separating Noether number of the non-abelian semidirect product $\mathrm{C}_{p}\rtimes\mathrm{C}_{3}$ equals $p+1$, under the assumption that the base field $K$ contains an element of multiplicative order $3p$ ([9] works over an algebraically closed base field, but the representation providing the exact lower bound for the separating Noether number is defined over any field containing an element of multiplicative order $3p$, and in order to prove that $p+1$ is also an upper bound, we may pass to the algebraic closure using Lemma 3.1).
@@ 8.4. The group $\mathrm{M}_{27}$
Now (37), (38) clearly imply that $w_{2}$ is a non-zero scalar multiple of $[1,1,1]^{T}$, $[\varepsilon,\varepsilon^{2},1]^{T}$, or $[\varepsilon^{2},\varepsilon,1]^{T}\}$, implying in turn by Lemma 8.12 that $\mathrm{Stab}_{G}(w_{2})\in\{\langle b\rangle,\langle a^{3}b\rangle,\langle a^{6}b\rangle\}$.

### v2

@@ 6.1. The group $\mathrm{Dic}_{8}\times\mathrm{C}_{2}$
(note that $\langle a^{2}\rangle$ is the commutator subgroup of $G$, so $a^{2}$ is in the kernel of any $1$-dimensional representation of $G$). For $j=1,2,3,4$ denote by $W_{j}$ the vector space $K^{2}$ endowed with the representation $\psi_{j}$, and for $\chi\in\widehat{G}$ denote by $U_{\chi}$ the vector space $K$ endowed with the representation $\chi$, and set $U:=\bigoplus_{\chi\in\widehat{G}}U_{\chi}$.
@@ 7. The group $\mathrm{S}_{3}\times\mathrm{C}_{3}$
For example, $q_{xx}=x_{1}x_{2}$, $q_{yz}=\frac{1}{2}(y_{1}z_{2}+y_{2}z_{1})$, $p_{xxx}=x_{1}^{3}+x_{2}^{3}$, $p_{xxy}=x_{1}^{2}y_{1}+x_{2}^{2}y_{2}$. It is well known (see for example [33, Theorem 4.1]) that the elements $q_{\xi\eta}$, $p_{\xi\eta\zeta}$ minimally generate $K[W_{1}\oplus W_{2}\oplus W_{3}]^{\mathrm{S}_{3}}$, where $\mathrm{S}_{3}$ is identified with the subgroup $\langle a,b\rangle$ of $G$. Set
@@
Assume next that $|\mathrm{Stab}_{G}(w)|=2$; this means that $H:=\mathrm{Stab}_{G}(w)$ is $\langle b\rangle$, $\langle ab\rangle$, or $\langle a^{2}b\rangle$, and $q_{yy}(w)\neq 0$. For any $H$-invariant monomial $T:=t_{\chi}\cdots t_{\chi^{\prime}}$ with $\deg(T)\leq 2$, either $T$, $q_{yy}T$, or $q_{yy}^{2}T$ is a $G$-invariant of degree at most $6$. It follows that $T(u)=T(u^{\prime})$ for any $H$-invariant monomial in $K[U]$ of degree at most $2$. Therefore there exists an $h\in H$ with $h\cdot u=u^{\prime}$ (note that $\beta_{\mathrm{sep}}^{K}(H)=\beta^{K}(H)=2$).
Assume next that $|\mathrm{Stab}_{G}(w)|=3$; this means that $H:=\mathrm{Stab}_{G}(w)$ is $\langle ac\rangle$ or $\langle ac^{2}\rangle$, and $p^{-}_{yyy}(w)\neq 0$. For any $H$-invariant monomial $T:=t_{\chi}\cdots t_{\chi^{\prime}}$ with $\deg(T)\leq 3$, either $T$ or $p^{-}_{yyy}T$ is a $G$-invariant of degree at most $6$. It follows that $T(u)=T(u^{\prime})$ for any $H$-invariant monomial in $K[U]$ of degree at most $3$. Therefore there exists an $h\in H$ with $h\cdot u=u^{\prime}$ (note that $\beta_{\mathrm{sep}}^{K}(H)=\beta^{K}(H)=3$).
@@ 8.3. The group $\mathrm{C}_{7}\rtimes\mathrm{C}_{3}$
It is proved in [9, Theorem 4.1] that for any prime $p$ congruent to $1$ modulo $3$, the separating Noether number of the non-abelian semidirect product $\mathrm{C}_{p}\rtimes\mathrm{C}_{3}$ equals $p+1$, under the assumption that the base field $K$ contains an element of multiplicative order $3p$. Actually, [9] works over an algebraically closed base field, but the representation providing the exact lower bound for the separating Noether number is defined over any field containing an element of multiplicative order $3p$, and in order to prove that $p+1$ is also an upper bound, we may pass to the algebraic closure using Lemma 3.1.
@@ 8.4. The group $\mathrm{M}_{27}$
Now (38), (39) clearly imply that $w_{2}$ is a non-zero scalar multiple of $[1,1,1]^{T}$, $[\varepsilon,\varepsilon^{2},1]^{T}$, or $[\varepsilon^{2},\varepsilon,1]^{T}$, implying in turn by Lemma 8.12 that $\mathrm{Stab}_{G}(w_{2})\in\{\langle b\rangle,\langle a^{3}b\rangle,\langle a^{6}b\rangle\}$.
@@
Acknowledgements
This research was partially supported by the Hungarian National Research, Development and Innovation Office, NKFIH K 138828.